OpenStudy (anonymous):

OK mymathlab sucks, lol. Please help me figure this out... Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx+b. (4,9); 2x+y=4 the answer I got is -2/1x+21

OpenStudy (lalaly):

\[y=-2x+4\]the slope of the perpindicular line is the negative reciprocal of the slope of the given line slope of perpindicular line =\[\frac{1}{2}\] now\[y-y_1=m(x-x_1)\] \[y-9=\frac{1}{2} (x-4)\]

OpenStudy (lalaly):

\[y=\frac{1}{2}x+7\]

OpenStudy (anonymous):

it would be a -1/2 right? the perpendicular line

OpenStudy (lalaly):

no coz its already -2 so the negative reciprocal of -2 is 1/2

OpenStudy (anonymous):

the original was a 2x I had figured the -2 as my answer.

OpenStudy (lalaly):

it has to be in the form y=mx+b

OpenStudy (anonymous):

ok I got ya, now I see my mistake, thank you so much.

OpenStudy (lalaly):

no problem=)